# The Non Parametric Reconstruction of Binary Black Hole Mass Evolution from GWTC-4.0 Gravitational Wave Catalog

### v1

@@ 2.3 Expected Number of Detections
$N_{\mathrm{exp}}(m_{k},z_{j}|\boldsymbol{\theta})$ $=T_{\rm obs}\int_{m_{k}}^{m_{k}+\Delta m_{k}}\int_{z_{j}}^{z_{j}+\Delta z_{j}}\mathcal{R}(z)\frac{p(m|z,\boldsymbol{\theta})}{1+z}$
$\times\frac{dV_{c}}{dz}S(m,z)\,dm\,dz,$ (6)
where $T_{\rm obs}$ is the total observation time, $\mathcal{R}(z)$ is the merger rate density per unit comoving volume at redshift $z$, $p(m|z,\boldsymbol{\theta})$ is the normalized mass distribution at redshift $z$, $(1+z)^{1}$ accounts for cosmological time dilation, and $dV_{c}/dz=4\pi cd_{L}^{2}(z)/[(1+z)^{2}H(z)]$ is the differential comoving volume per unit redshift.
For computational efficiency, we factorize this expression. Assuming the merger rate evolution can be absorbed into our Taylor expansion framework, we write:
$N_{\mathrm{exp}}(m_{k},z_{j}|\boldsymbol{\theta})$ $=N_{\mathrm{tot}}\cdot\tilde{p}(m_{k},z_{j}|\boldsymbol{\theta})\cdot S(m_{k},z_{j})$
$\times\frac{dV_{c}}{dz}\bigg|_{z_{j}}\cdot\frac{1}{1+z_{j}}\cdot\Delta m_{k}\Delta z_{j},$ (7)
where $N_{\mathrm{tot}}=\sum_{k,j}N_{\mathrm{obs}}(m_{k},z_{j})$ is the total number of observed events, and $\tilde{p}(m_{k},z_{j}|\boldsymbol{\theta})$ is our Taylor-expanded mass-redshift distribution. The factor $dV_{c}/dz|_{z_{j}}$ provides the comoving volume element at each redshift bin, naturally weighting the contribution of each bin by the available survey volume. The factor $(1+z_{j})^{-1}$ accounts for the cosmological time dilation.
2.4 Likelihood and Inference
The population parameters are $\boldsymbol{\theta}=\{p_{0}(m_{k}),p_{1}(m_{k}),p_{2}(m_{k})\}_{k=1}^{N_{m}}$. We require the intrinsic model to remain positive and properly normalized across all bins.
Assuming Poisson statistics for the observed counts in each bin, the likelihood is:
$\mathcal{L}(\boldsymbol{\theta})=\prod_{k,j}\frac{N_{\mathrm{exp}}(m_{k},z_{j}|\boldsymbol{\theta})^{N_{\mathrm{obs}}(m_{k},z_{j})}e^{N_{\mathrm{exp}}(m_{k},z_{j}|\boldsymbol{\theta})}}{N_{\mathrm{obs}}(m_{k},z_{j})!}.$ (8)
This Poisson likelihood models the observed counts in each bin denoted by the index k, and j, where $N_{\mathrm{exp}}(m_{k},z_{j})$ represents the expected number of detections given the intrinsic distribution and selection effects, while $N_{\mathrm{obs}}(m_{k},z_{j})$ are the actual observed counts. The Poisson distribution naturally accounts for the discrete nature of GW detections and the statistical fluctuations in sparse bins, making it particularly suitable for the relatively small sample sizes typical of current GW catalogs. To account for parameter estimation uncertainties, we use the full posterior sample of every event, ensuring full posterior support is explored across the ensemble. Events spanning multiple bins naturally contribute to coefficient uncertainties through this sampling procedure.
We adopt priors that ensure physical behavior while remaining minimally informative. We use uniform priors $p_{1}(m_{k})\sim\text{Uniform}(-4,4)$ and $p_{2}(m_{k})\sim\text{Uniform}(-4,4)$ that encompass physically reasonable evolutionary timescales. We sample from this posterior using ensemble Monte Carlo methods.We sample from the posterior using the emcee ensemble sampler (Foreman-Mackey et al., 2013). This framework thus provides a model independent probe of BBH mass function evolution, with coefficients that have direct astrophysical interpretations.
3 Observed BBH Population and Evidence for Mass Distance Trends
Figure 2: Exploratory visualization of observed BBH component mass and distance distributions from the GWTC-3 and GWTC-4.0 catalogs. The top left panel shows a scatter plot of median posterior component masses versus median luminosity distance for GWTC-3, categorized as low (20-40 M, blue), medium (40-60 M, purple), and high (>>60 M, orange) mass ranges, highlighting the distribution of detected systems. The top right panel presents violin plots for GWTC-3, with the left violins showing the distributions of primary component masses and the right violins showing distributions for secondary masses, grouped in three luminosity distance bins (0-2, 2-4.5, and 4.5-8 Gpc), illustrating how both mass components vary with distance. The bottom left panel repeats the mass versus luminosity distance scatter plot for GWTC-4.0, revealing analogous mass-dependent distance patterns. The bottom right panel displays violin plots for GWTC-4.0, again distinguishing primary (left) and secondary (right) mass distributions within each distance bin, revealing the changing spread and range of observed masses at increasing distances. Collectively, these panels illustrate the evolving mass-distance relationship across observing runs and motivate the application of hierarchical population analyses that rigorously account for uncertainties and selection biases. Importantly, low-mass BBHs are not detectable at higher redshifts primarily due to selection effects, while the absence of high-mass BBHs at lower distances cannot be explained by selection biases alone, suggesting intrinsic astrophysical factors influence their local detectability.
@@ 3 Observed BBH Population and Evidence for Mass Distance Trends
Figure 2 presents event-level summary statistics illustrating qualitative trends in the detected samples from the GWTC-3 and GWTC-4.0 catalogs. The top row shows results for GWTC-3, with the left panel displaying median component mass versus median luminosity distance for each posterior, grouped into three mass categories (low: 20-40 M, medium: 40-60 M, high: >>60 M). The right panel shows violin plots of primary (left) and secondary (right) component mass distributions within each luminosity distance bin (0-2, 2-4.5, and 4.5-8 Gpc), demonstrating how both mass components vary with distance. The bottom row provides the corresponding GWTC-4.0 results, with the left panel depicting the mass versus distance scatter plot and the right panel presenting violin plots for primary (left) and secondary (right) mass within the same luminosity distance bins. Together, these visual summaries reveal a stronger correlation between component mass and distance in GWTC-4.0 compared to GWTC-3, consistent with GWTC-4.0’s broader redshift coverage. These trends confirm that higher-mass BBHs tend to be detected at greater distances (impacted by selection effects) and exhibit broader mass distributions across distance ranges.
@@
Such patterns can arise from two non exclusive effects. First, heavier binaries generate stronger GW signals and are therefore detectable at larger distances; this selection bias alone can produce an apparent mass redshift correlation even if the intrinsic population is non evolving. Second, there may be genuine astrophysical evolution of the BBH mass function for example, metallicity driven differences in stellar winds and mass loss, that enhances the formation of massive remnants at earlier cosmic times. Disentangling these two effects requires quantitative correction for detector sensitivity and measurement uncertainty.
@@
Figure 3: Posterior constraints on the redshift evolution of the BBH mass distribution, shown as a function of component mass. Left: Linear coefficient $p_{1}(m)$, measuring the first-order rate of change of the mass distribution with redshift. Right: Quadratic coefficient $p_{2}(m)$, capturing any curvature in the redshift dependence. Error bars denote $1\sigma$ credible intervals from the hierarchical Bayesian analysis. Two binning schemes are presented: a coarse $5\times 5$ grid with $\Delta m=10\,M_{\odot}$ and $\Delta z=0.2$, and a finer $10\times 10$ grid with $\Delta m=5\,M_{\odot}$ and $\Delta z=0.1$. The latter exhibits larger uncertainties consistent with reduced event counts per bin. The linear term is consistent with zero for low mass systems ($m\lesssim 30\,M_{\odot}$) but increases toward higher masses, indicating massive black holes are relatively more abundant at earlier cosmic times. The quadratic term remains consistent with zero across all mass bins, implying a simple linear dependence on redshift adequately describes the current observational range $z\lesssim 1$ accessible to LVK.
@@
To distinguish between genuine evolution of the BBH mass distribution and observational biases, we employ the LVK injection campaigns Observing Runs O1-O4. These datasets contain simulated signals with known source parameters, analyzed through the same search pipelines as the real data, thereby providing an empirical estimate of detection efficiency across the $(m,z)$ plane. For consistency with the observed catalog, we adopt a conservative network signal to noise threshold of $\rho_{\rm thresh}>10$ when defining detections in both the injections and the real sample. The resulting selection function, together with the full posterior samples of the observed events, is incorporated into our hierarchical Bayesian framework (Section 2). This framework models the intrinsic joint distribution $p(m,z)$ via a non parametric Taylor expansion, while explicitly marginalizing over measurement uncertainties and correcting for selection effects. The results of this analysis are presented in Section 4.
It is important to emphasize that Fig. 2 is exploratory only. The plotted medians and means provide intuition about the detected sample but do not constitute a statistically rigorous analysis, as they neglect selection effects and the full posterior structure of individual events.
4 Results
@@
We apply the non parametric Bayesian framework described in Section 2 to the GWTC-3 and GWTC-4.0 catalog of BBH mergers, incorporating the corresponding injection campaigns to account for detector selection effects. The analysis is performed in the $(m,z)$ plane, discretized into grids of varying resolutions to study the impact of bin size on the results. We consider two cases: one with a $5\times 5$ grid characterized by $\Delta m=10\,M_{\odot}$ and $\Delta z=0.2$, and another with a finer $10\times 10$ grid having $\Delta m=5\,M_{\odot}$ and $\Delta z=0.1$. Both cases cover the redshift range $0\leq z\leq 1$ and mass range $30\,M_{\odot}\leq m\leq 80\,M_{\odot}$.
To ensure robustness, we restrict the analysis to systems with component masses above $30\,M_{\odot}$. At lower masses the GWTC-3 and GWTC-4.0 sample and the injection sets exhibit rapidly decreasing detection efficiency, sparse injection coverage, and larger relative mass uncertainties, which together render the inverse selection correction unstable. Limiting to $m\gtrsim 30\,M_{\odot}$ ensures that the inferred Taylor coefficients are driven by regions of the $(m,z)$ plane where both the observed events and the injections provide reliable information.
The detector selection function $S_{(}m,z)$ is calibrated from the injection campaign of LVK collaboration. For each bin $(m_{k},z_{j})$, we compute the detection probability as the fraction of injected signals recovered above the network SNR threshold (see Sec. 2), and then apply volume–time corrections to account for cosmological effects. The resulting selection function exhibits the expected strong dependence on both mass and redshift (Gerosa & Bellotti, 2024; Essick et al., 2025).
Figure 3 summarizes the key results of our non-parametric inference. The panels show the marginalized posterior distributions of the linear ($p_{1}$) and quadratic ($p_{2}$) evolution coefficients as functions of component mass. A clear trend emerges in the linear term: for lower mass systems ($m\lesssim 30\,M_{\odot}$), both $p_{1}(m)$ and $p_{2}(m)$ are consistent with zero, indicating no detectable redshift evolution. At higher masses, however, $p_{1}(m)$ shifts systematically toward positive values, suggesting that more massive black holes were relatively more common at earlier cosmic times. This pattern aligns with astrophysical models where metallicity influences the efficiency of forming massive remnants. Conversely, the quadratic coefficient $p_{2}(m)$ shows no statistically significant deviation from zero across all mass bins, implying that a simple linear and quadratic dependence on redshift sufficiently describes the currently accessible range ($z\lesssim 1$) probed by current LVK observations. It is important to also note that due to large number of bins present in this non-parametric model independent approach, the statistical uncertainties on the reconstruction is shadowing to draw any conclusive measurement of the redshift dependent mass distribution. A well motivated theoretical study with fewer number of parameters can unveil any potential trend with a better statistical significance.
These constraints are presented for two discretizations schemes in the $(m,z)$ plane: a coarser $5\times 5$ grid with $\Delta m=10\,M_{\odot}$, $\Delta z=0.2$ and a finer $10\times 10$ grid with $\Delta m=5\,M_{\odot}$, $\Delta z=0.1$. The finer binning reveals more detailed features in the evolution coefficients but comes with larger statistical uncertainties, as reflected in the increased error bars. This comparison highlights the trade-off between resolution and statistical significance when studying BBH mass evolution over redshift.
@@
5 Conclusions
We have developed and applied a fully non-parametric Bayesian framework to infer the redshift evolution of the BBH mass distribution using all confident BBH detections from the LVK GWTC-3 and GWTC-4.0 catalogs.
@@ 5 Conclusions
By expanding the intrinsic joint distribution $p(m,z)$ as a Taylor series around $z_{\rm ref}=0.25$, the method captures evolutionary trends without imposing restrictive parametric assumptions, while rigorously accounting for detector selection effects through injection calibrated efficiencies.
@@
Our analysis yields tentative but consistent evidence for mass dependent evolution. The linear evolution coefficient $p_{1}(m)$ is consistent with zero for lower mass systems ($m\lesssim 30\,M_{\odot}$), but becomes increasingly positive at higher masses. In particular, the highest mass bins show tentative evidence that massive black holes were relatively more abundant at earlier cosmic times. By contrast, the quadratic term $p_{2}(m)$ remains consistent with zero across all mass ranges, indicating that a linear description suffices up to the redshift range $z\sim 1$ currently probed by LVK observations.
@@
These results are qualitatively consistent with theoretical expectations for metallicity driven evolution: low metallicity environments at high redshift suppress stellar winds, allowing the formation of more massive stellar cores and correspondingly heavier black hole remnants. The lack of detectable evolution for lower mass systems suggests that their formation channels are less sensitive to environmental variations, consistent with robust core collapse pathways. The approximately linear trend indicates gradual, metallicity driven changes over cosmic time rather than abrupt evolutionary transitions. This result agrees with the probability of different formation channels, we found in GWTC-4.0 using phase-space, showing that above approximately $45$ M there is an indication of a different population of BBHs (Afroz & Mukherjee, 2025c).
@@
Beyond these specific findings, this work demonstrates the power of non parametric approaches for GW population inference. Unlike parametric models, which may obscure or bias subtle features, our Taylor expansion framework provides a flexible yet interpretable means of capturing population trends while fully correcting for observational selection effects (see Section 2). The method’s ability to reveal tentative mass dependent evolution highlights its value for future analyses with larger datasets.
@@
Previous studies based on GWTC-3 (Fishbach et al., 2021; van Son et al., 2022; Tiwari, 2022; Karathanasis et al., 2023; Rinaldi et al., 2024; Heinzel et al., 2025b; Lalleman et al., 2025) reported mixed evidence for mass–redshift evolution, ranging from no detectable trends to weak indications at high masses. By incorporating the larger GWTC-4.0 along with GWTC-3 catalog, our analysis reveals tentative evidence for mass-dependent linear evolution.

### v2

@@ II.3 Hierarchical Inference Framework
$\rm{\mathcal{L}(\boldsymbol{\theta}_{kj})=e^{-N_{\rm exp}(m_{k},z_{j}\mid\boldsymbol{\theta}_{kj})}\prod_{i=1}^{N_{\rm obs}}\left[W_{ijk}\,\frac{N_{\rm int}(m_{k},z_{j}\mid\boldsymbol{\theta}_{kj})}{\Delta m_{k}\,\Delta z_{j}}\right]}.$ (8)
here, the product is over the observed events.
III Observed BBH Population and Evidence for Mass Distance Trends
Figure 1: Mass-redshift distributions for the primary (left) and secondary (right) black hole components in GWTC-4 catalog. Each distribution combines individual-event posterior samples with prior weighting.
@@ III Observed BBH Population and Evidence for Mass Distance Trends
Figure 1 presents the resulting mass-redshift distributions of primary mass (left) and secondary mass (right) constructed from these posterior samples across all 176 events. Each 2D histogram uses prior reweighting to recover the intrinsic population-level distribution, revealing the characteristic BBH mass hierarchy and redshift evolution over $\rm{m\in[5,100]}\,$M, $\rm{z\in[0,1.2]}$.
@@
The observed distribution of the GW sources in the mass and redshift plane can arise from two non exclusive effects. First, heavier binaries generate stronger GW signals and are therefore detectable at larger distances; this selection bias alone can produce an apparent mass redshift correlation even if the intrinsic population is non evolving. Second, there may be genuine astrophysical evolution of the BBH mass function for example, metallicity driven differences in stellar winds and mass loss, that enhances the formation of massive remnants at earlier cosmic times. Disentangling these two effects requires quantitative correction for detector sensitivity and measurement uncertainty.
@@
Figure 2: Posterior constraints on the redshift evolution of the BBH primary mass ($\rm{m_{1}}$) distribution for a binning scheme with $\rm{\Delta M=10\,M_{\odot}}$ and $\rm{\Delta z=0.1}$. Left: Constant coefficient $\rm{\log(P_{0})}$ describing the baseline primary mass distribution. Middle: Linear coefficient $\rm{P_{1}}$, quantifying the first-order variation of the mass distribution with redshift. Right: Quadratic coefficient $\rm{P_{2}}$, probing possible curvature in the redshift dependence. Error bars represent $\rm{1\sigma}$ credible intervals obtained from the hierarchical Bayesian analysis. The linear coefficient is consistent with zero at low masses ($\rm{m_{1}\lesssim 30\,M_{\odot}}$) and shows a mild positive trend toward higher masses, suggesting that more massive black holes may be relatively more abundant at earlier cosmic times. The quadratic coefficient remains consistent with zero across all mass bins, indicating that a linear redshift dependence adequately describes the population within the current observational range $\rm{z\lesssim 1}$ accessible to LVK.
Figure 3: Posterior constraints on the redshift evolution of the BBH secondary mass ($\rm{m_{2}}$) distribution for a binning scheme with $\rm{\Delta M=10\,M_{\odot}}$ and $\rm{\Delta z=0.1}$. Left: Constant coefficient $\rm{\log(P_{0})}$ describing the baseline primary mass distribution. Middle: Linear coefficient $\rm{P_{1}}$, quantifying the first-order variation of the mass distribution with redshift. Right: Quadratic coefficient $\rm{P_{2}}$, probing possible curvature in the redshift dependence. Error bars represent $\rm{1\sigma}$ credible intervals obtained from the hierarchical Bayesian analysis. The linear coefficient is consistent with zero at low masses ($\rm{m_{1}\lesssim 30\,M_{\odot}}$) and shows a mild positive trend toward higher masses, suggesting that more massive black holes may be relatively more abundant at earlier cosmic times. The quadratic coefficient remains consistent with zero across all mass bins, indicating that a linear redshift dependence adequately describes the population within the current observational range $\rm{z\lesssim 1}$ accessible to LVK.
Figure 4: Posterior constraints on the redshift evolution of the BBH primary mass ($\rm{m_{1}}$) distribution for a binning scheme with $\rm{\Delta M=5\,M_{\odot}}$ and $\rm{\Delta z=0.1}$. Left: Constant coefficient $\rm{\log(P_{0})}$ describing the baseline primary mass distribution. Middle: Linear coefficient $\rm{P_{1}}$, quantifying the first-order variation of the mass distribution with redshift. Right: Quadratic coefficient $\rm{P_{2}}$, probing possible curvature in the redshift dependence. Error bars represent $\rm{1\sigma}$ credible intervals obtained from the hierarchical Bayesian analysis. The linear coefficient is consistent with zero at low masses ($\rm{m_{1}\lesssim 30\,M_{\odot}}$) and shows a mild positive trend toward higher masses, suggesting that more massive black holes may be relatively more abundant at earlier cosmic times. The quadratic coefficient remains consistent with zero across all mass bins, indicating that a linear redshift dependence adequately describes the population within the current observational range $\rm{z\lesssim 1}$ accessible to LVK.
Figure 5: Posterior constraints on the redshift evolution of the BBH secondary mass ($\rm{m_{2}}$) distribution for a binning scheme with $\rm{\Delta M=5\,M_{\odot}}$ and $\rm{\Delta z=0.1}$. Left: Constant coefficient $\rm{\log(P_{0})}$ describing the baseline primary mass distribution. Middle: Linear coefficient $\rm{P_{1}}$, quantifying the first-order variation of the mass distribution with redshift. Right: Quadratic coefficient $\rm{P_{2}}$, probing possible curvature in the redshift dependence. Error bars represent $\rm{1\sigma}$ credible intervals obtained from the hierarchical Bayesian analysis. The linear coefficient is consistent with zero at low masses ($\rm{m_{1}\lesssim 30\,M_{\odot}}$) and shows a mild positive trend toward higher masses, suggesting that more massive black holes may be relatively more abundant at earlier cosmic times. The quadratic coefficient remains consistent with zero across all mass bins, indicating that a linear redshift dependence adequately describes the population within the current observational range $\rm{z\lesssim 1}$ accessible to LVK.
@@
To distinguish between genuine evolution of the BBH mass distribution and observational biases, we employ the LVK injection campaigns Observing Runs O1-O4. The selection function, together with the full posterior samples of the observed events, is incorporated into our hierarchical Bayesian framework (Section II). This framework models the intrinsic joint distribution $\rm{p(m,z)}$ via Taylor expansion, while explicitly marginalizing over measurement uncertainties and correcting for selection effects. The results of this analysis are presented in Section IV.
IV Results
@@
We apply the non parametric Bayesian framework described in Section II to the GWTC-3 and GWTC-4 catalog of BBH mergers, incorporating the corresponding injection campaigns to account for detector selection effects. The analysis is performed in the $\rm{(m,z)}$ plane, discretized into grids of varying resolutions to study the impact of bin size on the results. We consider two cases: one with $\rm{\Delta m=10\,M_{\odot}}$ and $\rm{\Delta z=0.1}$, and another with a finer grid having $\rm{\Delta m=5\,M_{\odot}}$ and $\rm{\Delta z=0.1}$. Both cases cover the redshift range $\rm{0\leq z\leq 1}$ and mass range $\rm{10\,M_{\odot}\leq m\leq 80\,M_{\odot}}$. The detector selection function $\rm{S(m,z)}$ is calibrated from the injection campaign of LVK collaboration.
The Bayesian inference was performed using the emcee (Foreman-Mackey et al., 2013) affine-invariant ensemble Markov Chain Monte Carlo (MCMC) sampler. For the finer mass binning configuration ($\rm{\Delta M=5\,M_{\odot}}$), we used $200$ walkers and ran the chains for 50 00050\,000 steps, discarding the first 10 00010\,000 steps as burn-in. For the coarser binning configuration ($\rm{\Delta M=10\,M_{\odot}}$), we used $100$ walkers with 25 00025\,000 steps and removed the first 5 0005\,000 steps as burn-in. Uniform priors were adopted for all Taylor expansion coefficients describing the mass evolution, with $\rm{\log P_{0}\in[-13,0]}$, $\rm{P_{1}\in[-20,20]}$, and $\rm{P_{2}\in[-30,30]}$.
Figures 2-5 summarize the key results of our non-parametric reconstruction of the redshift evolution of the BBH mass distribution. We present constraints separately for the primary mass ($\rm{m_{1}}$) and secondary mass ($\rm{m_{2}}$). For the primary mass ($\rm{m_{1}}$), a clear qualitative trend emerges in the linear coefficient $\rm{p_{1}(m_{1})}$. In both binning schemes, $\rm{p_{1}(m_{1})}$ is consistent with zero at low masses ($\rm{m_{1}\lesssim 25\text{-}30,M_{\odot}}$), indicating no statistically significant redshift evolution in this regime. Toward higher masses, however, $\rm{p_{1}(m_{1})}$ shifts systematically toward positive values. In the large-bin case, this manifests as a smooth rise of the linear coefficient beyond $\rm{\sim 40,M_{\odot}}$, while in the finer binning the trend becomes more structured but remains qualitatively similar. This behavior suggests that more massive primary black holes were relatively more abundant at earlier cosmic times, consistent with astrophysical scenarios in which lower metallicity environments at higher redshift favor the formation of heavier remnants. The full posterior distributions of the inferred parameters, including the corresponding corner plots for both binning configurations, are presented in Appendix A.
In contrast, the quadratic coefficient $\rm{p_{2}(m_{1})}$ remains statistically consistent with zero across nearly all mass bins for both discretizations. Although fluctuations are visible, particularly in the finer binning, they are well within the $\rm{1\sigma}$ uncertainties. This indicates no compelling evidence for strong curvature in the redshift dependence of the primary mass distribution within the currently accessible range $\rm{z\lesssim 1}$ probed by LVK.
A broadly similar picture emerges for the secondary mass ($\rm{m_{2}}$), albeit with larger statistical uncertainties. For the coarser binning ($\rm{\Delta m=10,M_{\odot}}$), the linear coefficient $\rm{p_{1}(m_{2})}$ shows a gradual increase toward higher masses, with low-mass bins consistent with zero and high-mass bins favoring positive values. In the finer binning, this increasing trend persists but appears noisier, with larger error bars reflecting the reduced number of events per bin. Nevertheless, the overall pattern is consistent with that observed for the primary mass: a tendency for more massive components to exhibit stronger positive redshift evolution.
The quadratic coefficient for secondary mass $\rm{p_{2}(m_{2})}$ does not display any statistically significant deviation from zero across the mass range. While individual bins fluctuate above or below zero, the uncertainties remain large and encompass the null hypothesis of no quadratic evolution. Therefore, within current observational precision, a linear redshift dependence appears sufficient to describe the secondary mass evolution as well.
Our analysis marginalizes over mass ratio and spin distributions by working directly with component masses and using selection functions from comprehensive injection campaigns. Black hole spins remain poorly constrained in current GW observations, and we do not model their potential redshift evolution. However, the mass-dependent detector sensitivity is dominated by the chirp mass dependence of GW amplitude, with spins contributing subdominant corrections, ensuring our mass evolution inferences are robust to spin distribution uncertainties.
Our findings can be compared with previous population studies using GWTC-3. Fishbach et al. (2021) found hints of increasing characteristic mass with redshift using parametric models, qualitatively consistent with our mildly positive $\rm{p_{1}(m)}$ at high masses. Tiwari (2022) reported results consistent with no evolution across all masses, which aligns with our findings at $\rm{m\lesssim 30\,M_{\odot}}$ but differs at higher masses where very weak indications of evolution. Karathanasis et al. (2023) explored metallicity-dependent evolution scenarios and found weak evidence for mass evolution, corroborating our mass-dependent trends. The inclusion of GWTC-4 data, with its extended redshift reach and increased event count, allows our non-parametric approach to reveal mass-resolved evolutionary features that may have been obscured in earlier parametric analyses or smaller catalogs.
@@
V Conclusions
@@
We have developed and applied a fully non-parametric Bayesian framework to infer the redshift evolution of the BBH mass distribution using all confident BBH detections from the LVK GWTC-3 and GWTC-4 catalogs. By expanding the intrinsic joint distribution $\rm{p(m,z)}$ as a Taylor series around $\rm{z_{\rm ref}=0.25}$, the method captures evolutionary trends without imposing restrictive parametric assumptions, while rigorously accounting for detector selection effects through injection calibrated efficiencies.
@@
Our analysis yields very mild evidence for mass-dependent evolution. The linear evolution coefficient $\rm{p_{1}(m)}$ is consistent with zero for lower mass systems ($\rm{m\lesssim 30\,M_{\odot}}$), but becomes increasingly positive at higher masses. In particular, the highest mass bins show very mild indications that massive black holes may have been relatively more abundant at earlier cosmic times. By contrast, the quadratic term $\rm{p_{2}(m)}$ remains consistent with zero across all mass ranges, indicating that a linear description suffices up to the redshift range $\rm{z\sim 1}$ currently probed by LVK observations.
@@
These results are qualitatively consistent with theoretical expectations for metallicity driven evolution: low metallicity environments at high redshift suppress stellar winds, allowing the formation of more massive stellar cores and correspondingly heavier black hole remnants. The lack of detectable evolution for lower mass systems suggests that their formation channels are less sensitive to environmental variations, consistent with robust core collapse pathways. The approximately linear trend indicates gradual, metallicity driven changes over cosmic time rather than abrupt evolutionary transitions. This trend is qualitatively consistent with the probability of different formation channels inferred from phase-space analyses, showing that above approximately $45$ M there is an indication of a different population of BBHs (Afroz & Mukherjee, 2025c).
@@
Beyond these specific findings, this work demonstrates the power of non-parametric approaches for GW population inference. Unlike parametric models, which may obscure or bias subtle features, our Taylor expansion framework provides a flexible yet interpretable means of capturing population trends while fully correcting for observational selection effects (see Section II). The method’s ability to probe possible mass-dependent evolution highlights its value for future analyses with larger datasets.
@@
Previous studies based on GWTC-3 (Fishbach et al., 2021; van Son et al., 2022; Tiwari, 2022; Karathanasis et al., 2023; Rinaldi et al., 2024; Heinzel et al., 2025b; Lalleman et al., 2025) reported mixed evidence for mass-redshift evolution, ranging from no detectable trends to weak indications at high masses. Some works found hints of increasing characteristic mass with redshift (Fishbach et al., 2021), while others reported results consistent with no evolution (Tiwari, 2022). Our combined GWTC-3 and GWTC-4 analysis, employing a complementary non-parametric approach with mass resolved evolution coefficients, finds very mild evidence for evolution at high masses, while providing new insights into where (in mass) such evolution may be most pronounced.